OpenStudy (anonymous):

Let f(x) = 4 - 12/x^2. Calculate the lower and upper sum estimates of the area under the curve over the interval [2,5] using subintervals of width ½. I tried working it out but I screw up because I get a negative number.

OpenStudy (lgbasallote):

what do you mean by lower and upper sums? im not familiar with this fancy terminology..

OpenStudy (anonymous):

Maybe my textbook is weird

OpenStudy (turingtest):

right and left hand sums

OpenStudy (anonymous):

So wait are these Riemann sums o.o

OpenStudy (turingtest):

yeah it seems so look under methods to see what I think they mean by under and over or whatever... http://en.wikipedia.org/wiki/Riemann_sum

OpenStudy (turingtest):

these problems take a good long time to do...

OpenStudy (lgbasallote):

@TuringTest is it just \[\huge \int_2^5 (4 - \frac{12}{x^2})dx\]

OpenStudy (lgbasallote):

or no?

OpenStudy (turingtest):

that is exact, we want approximations, from which we can derive that formula

OpenStudy (turingtest):

that is the limit as the number of rectangles goes to infinity @lgbasallote

OpenStudy (anonymous):

I have a feeling this is Riemann but I wasn't integrating just simplifying the summations and plugging in the value for n within the interval

OpenStudy (anonymous):

my brain hurts

OpenStudy (turingtest):

yes, it will just be a summation, which is usually more of a pain than integration

OpenStudy (turingtest):

first let's look at the definition: the area under a curve f(x) is approximated by\[A\approx\sum_{i=1}^nf(x_i^*)\Delta x\]

OpenStudy (turingtest):

n is the number of rectangles we are going to use and \(\Delta x\) is the width of those rectangles

OpenStudy (turingtest):

you should be able to figure out \(\Delta x\) at this point

OpenStudy (turingtest):

oh it's given here then from that you should be able to figure out n, the number of rectangles we will use

OpenStudy (anonymous):

6, because the interval is [2,5] and the rectangles are 1/2 width.. yeah I have work for all of this already but I went wrong somewhere with the number crunching... the concept is just generally muddled confusion for me right now

OpenStudy (turingtest):

so the left hand endpoints are going to be from 2 to 4.5 and the rh ones are from 2.5 to 5

OpenStudy (turingtest):

left sum\[A\approx\sum_{i=0}^nf(x_i^*)\Delta x\]right sum\[A\approx\sum_{i=1}^nf(x_i^*)\Delta x\]and since \(\Delta x=\frac12\) and the starting point is \(a=2\) we have \[x_i^*=a+i\Delta x=2+\frac i2\]so we need to do two sums

OpenStudy (anonymous):

...Crap, hang on, I'm checking my textbook and Riemann is in the next section

OpenStudy (turingtest):

typo* the left hand sum is\[A\approx\sum_{i=0}^{n-1}f(x_i^*)\Delta x\]the right hand sum is \[A\approx\sum_{i=1}^nf(x_i^*)\Delta x\]

OpenStudy (anonymous):

@TuringTest Here is my textbook example I don't think it's Riemann exactly, but can you work off of that?

OpenStudy (anonymous):

and we're solving for n =6

OpenStudy (anonymous):

What is the solution?

OpenStudy (anonymous):

@satellite73 I'm curious, what is the solution to this question?